# Strong coalitions in graphs

### v1

@@ 5 Strong coalition number for some classes of graphs
Subcase (iii). Let $V_{2}=\{v_{2},v_{5},v_{10}\}$, $V_{3}=\{v_{7}\}$, and $V_{4}=\{v_{8}\}$ such that $V_{2}\cup V_{3}$ and $V_{2}\cup V_{4}$ are only strong coalitions of order 4 of $P_{11}$. As before, each of $V_{5}$ and $V_{6}$ must be in a strong coalition with $V_{1}$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ have no four vertices in common. So, $SC(P_{11})\leq 5$.
@@
Subcase (iv). Assume that three of singleton sets form strong coalition with $V_{2}$, which is impossible, as the graph $P_{11}$ has only two strong coalition of order 4.
@@
Case 4. $|V_{1}|=4,|V_{2}|=|V_{3}|=2$, and $|V_{i}|=1$ for $4\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, no two singleton sets can produce a strong coalition. Furthermore, no singleton set forms a strong coalition with $V_{2}$ and $V_{3}$. So, each of $V_{j}$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$. By Theorem 3.6, $V_{1}$ has at most three strong coalition partners, a contradiction. Hence, $SC(P_{11})\leq 5$.
@@
Case 7. $|V_{i}|=2$ for $4\leq i\leq 6$ and $|V_{6}|=1$. We observe that no union of any two sets has cardinality 4. Consequently, there is no $SC(G)$-partition of order 6. Then, $SC(P_{11})\leq 5$.
@@
Finally, Let $n\geq 12$. By finding coalition partition of order 6 we have the result. We consider the following cases.
@@
Therefore we have the result. $\square$
@@ 6 Conclusion
This paper introduces the concept of the strong coalition in graphs and explores various properties related to its number. We have investigated the existence of strong coalition partitions and presented some lower bound and upper bounds for strong coalition number. We presented a sufficient condition for graphs $G$ of order at least $4$ with $SC(G)=|V(G)|$. Also, we have determined the precise values of ${SC}(P_{n})$, ${SC}(K_{r,s})$, and ${SC}(C_{n})$. We have outlined some unresolved problems and potential research directions related to the strong coalition number of graphs. However, there is still much work to be done in this area.

### v2

@@ 5 Strong coalition number for some classes of graphs
Subcase (iii). Let $V_{2}\cup V_{3}=\{v_{2},v_{4},v_{7},v_{10}\}$. Suppose that each of $V_{4}$, $V_{5}$, $V_{6}$ can be a strong coalition with $V_{1}$. It follows that $\bigcap_{k=4}^{6}|V_{1}\cup V_{k}|=4$, which is impossible, as the pairs $V_{1}\cup V_{4}$, $V_{1}\cup V_{5}$, and $V_{1}\cup V_{6}$ containing the remaining vertices have no four vertices in common. Thus, $SC(P_{11})\leq 5$.
Subcase (iv). Let $V_{2}=\{v_{2},v_{5},v_{10}\}$, $V_{3}=\{v_{7}\}$, and $V_{4}=\{v_{8}\}$ such that $V_{2}\cup V_{3}$ and $V_{2}\cup V_{4}$ are strong coalitions of order 4 of $P_{11}$. In addition, each of $V_{5}$ and $V_{6}$ can be in a strong coalition with $V_{1}$. Then $\bigcap_{k=5}^{6}|V_{1}\cup V_{k}|=4$, which is impossible, as the pairs $V_{1}\cup V_{5}$ and $V_{1}\cup V_{6}$ comprising the remaining vertices $v_{1},v_{3},v_{4},v_{6},v_{9}$, and $v_{11}$ have no four vertices in common. So, $SC(P_{11})\leq 5$.
Subcase (v). Assume that $V_{2}=\{v_{2},v_{7},v_{10}\}$, $V_{3}=\{v_{4}\}$, and $V_{4}=\{v_{5}\}$ such that $V_{2}\cup V_{3}$ and $V_{2}\cup V_{4}$ are strong coalitions of order 4 of $P_{11}$. As before, each of $V_{5}$ and $V_{6}$ can be in a strong coalition with $V_{1}$. It holds that $\bigcap_{k=5}^{6}|V_{1}\cup V_{k}|=4$, which is impossible, as the pairs $V_{1}\cup V_{5}$ and $V_{1}\cup V_{6}$ containing the remaining vertices have no four vertices in common. Thus, $SC(P_{11})\leq 5$.
@@
Subcase (vi). Without loss of generality, assume that each of $V_{3}$, $V_{4}$, and $V_{5}$ form a strong coalition with $V_{2}$. It is clear that $V_{6}$ can form a strong coalition with $V_{1}$. Then $\bigcap_{k=3}^{5}|V_{2}\cup V_{k}|=3$, which is impossible, as there are two sets such as $A$ and $C$ such that $|A\cap C|=2$.
@@
Case 4. $|V_{1}|=4,|V_{2}|=|V_{3}|=2$, and $|V_{i}|=1$ for $4\leq i\leq 6$. Since $\gamma_{st}(P_{11})=4$, no two singleton sets can produce a strong coalition. Furthermore, no singleton set forms a strong coalition with $V_{2}$ and $V_{3}$. So each of $V_{j}$ for $2\leq j\leq 6$ must be in a strong coalition with $V_{1}$, which contradicts Theorem 3.6. Hence, $SC(P_{11})\leq 5$.
@@
Case 7. $|V_{i}|=2$ for $1\leq i\leq 5$ and $|V_{6}|=1$. Then, $V_{6}$ together with any other set cannot have four vertices and hence cannot form a strong coalition. Consequently, there is no $SC(G)$-partition of order 6. Then, $SC(P_{11})\leq 5$.
@@
Finally, let $n\geq 12$. By Theorem 3.7, for any path $P_{n}$ we have $SC(P_{n})\leq 6$. Now, we find a maximal strong coalition partition of order 6 for $P_{n}$, where $n\geq 12$. Now, We consider the following cases.
@@
Therefore the proof is complete. $\square$
@@ 6 Conclusion
This paper introduces the concept of strong coalition in graphs. We have investigated the existence of strong coalition partitions and presented some lower bound and upper bounds for the strong coalition number. We presented a sufficient condition for graphs $G$ of order at least $4$ with $SC(G)=|V(G)|$. Also, we have determined the precise values of ${SC}(P_{n})$, ${SC}(K_{r,s})$, and ${SC}(C_{n})$. We have outlined some unresolved problems and potential research directions related to the strong coalition number of graphs. However, there is still much work to be done in this area.